Monthly Archives: March 2015

The week of March 23

We did some simulations of data transfer rates in preparation for VIRUS coming on line. The data set consists of 900 fits files; each file is 8524800 bytes in size. The working assumption is that one Hetdex image consists of 150 spectrographs x 2 amplifiers/spectrograph x 3 dither per field with a CCD size of 1032×2064 pixels. Total data size is 7,672,320,000 bytes (7.3 GB). The data would be collected over 20 minutes (~6 minutes per dither) but it only took 449.5 seconds (7.5 minutes) to transfer this data to Austin. Thus we should be able to keep up with the data collection rate without bringing the network to a crawl.

The large air and glycol (“Blue Thunder”) hoses were installed through the Az wrap and laid in the trench in the floor and through one of the ports through the pier to the ring wall in the last two weeks. The next step will be to connect these to the outside lines and do a pressure test of the system.

All power connections for the VIRUS enclosures are now complete and tested. We have also conducted some heat load tests on the liquid nitrogen system by placing plastic bags over each of the VIRUS bayonets (the component that will stick into a VIRUS unit to keep it cold). So far most of the system checks out close to specification.

The primary mirror group installed 4 mirrors this week and continue to keep up their 30% over expected swap rate in hopes of being close to completion when the corrector is ready for commissioning.

The CCAS shutter failed this week and could not be repaired from inside the CCAS Dome. Instead of having to rent a very expensive 90 ft manlift the mechanical team found a way to safely undo the mounting to the CCAS Dome and lower it to the ground. A replacement shutter is on order and should arrive in the next week or two. It will be installed by just reversing the procedure developed to remove the old shutter. Until then we’ve installed a simple wood plug that can be removed manually (after climbing the 90ft tower) to continue with critical primary mirror work.

The week of March 9

The grating and guard rail around the liquid nitrogen tank have been completed. This gives the contractors safer access to the controls during our filling.

The air and glycol (“Blue Thunder”) hoses arrived and installation may start as early as next week. This will actually shut down the structure for a day or two as these are carefully integrated into the azimuth cable wrap below the telescope.

The HRS instrumentation team came out from Austin and disassembled the HRS and packed the optics up. Some of the optics will just be stored while the coated optics are being shipped out to be recoated for the HRS upgrade. The HRS enclosure has been cleared and once some preliminary electrical work is done the inside of the enclosure will be painted flat black.

Our telescope operator and video production guru has completed a new video showing the installation of the temporary clean room frame. See This clean room will house the corrector when it arrives in late April.

A visiting engineer from Texas A&M was back this week with the doors and sides for the VIRUS enclosure. A lot of work continues to seal, plumb and install electrical components in the enclosures and annex.

In night time commissioning we attempted a rewind test. This is to measure the time it takes from the abort of one 20 minute trajectory and rewind and start a 2nd trajectory. It took 79 seconds to make this transition which is a little longer than we really wish but still within the 80 seconds from abort to guider acquisition. We hope to shave 9 seconds off this time so that we can have 10 seconds to have the metrology and guiders acquire their stars.

The mirror team has been getting ahead of their minimum swap rate by swapping another 5 mirrors this last week. They hope to have all of the older mirrors swapped out by the time we start re-commissioning the telescope.

The week of February 23

The last two weeks have not seen any obvious large changes to the facility. Some of the team continue to work on the VIRUS enclosures and their annex. The Annex is where much of the breakers, electronics, fans and junction boxes are located. Other members are sealing/caulking the seams of the enclosure so that when we close it up it will be air tight.

We did have some contractors on site finishing up a few punch list items from our remote thermal area project including a team from Trane came out last week to setup some of their communication hardware and test the glycol temperature control system as far as possible without a load on the system.

We painted some test squares of ultra flat black on the High Resolution Spectrograph (HRS) enclosure in preparation for the beginning of its upgrade.

The mirror team has swapped 4 mirror mirrors in the last two weeks keeping on schedule. They continue to make some improvements to the Strip and Wash room including installing a 2nd hot water hose in strip wash area and adding acid proof shelving.

The at angle tests with the wide field corrector at the University of Arizona are going well. So far all deflections detected are elastic and within specification. We are going to have a final review of the corrector alignment with several external experts in early April to get the final ok for shipping the corrector to the telescope.